Oct 9, 2014 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:30 a.m. Eastern (until switch from DST and then return to 11:00am)

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Add iron shields to top cosmic-ray counters (Elton)
  2. Check into procedure for new MOUs (Elton)
  3. Send Fcal plugin to drill down to block level energies (Matt to Will)

Tentative Agenda

  1. Announcements
  2. Action Items
  3. On call cell phone
  4. Preparations for beam [Checklists (Accessing Onsite Webpages From Offsite|offsite access)
  5. FCAL HDFCAL log book
  6. BCAL HDBCAL log book
  7. Reconstruction and Simulation
  8. Calibration
  9. Any other business


Attendees: Elton, Will, Mark (JLab); Matt (IU); Will L. (CMU); Tegan, Andrei, Noemi, Ahmed (UofR)

  1. Announcements
  2. Action Items
  3. On call cell phone
  4. Cross-training yesterday: FCAL, BCAL
    • Manuel and Elton went over the slow control settings for the two detectors. How to set and restore voltages.
    • We also went down to the hall and pointed out the location of the electronics.
  5. Preparations for beam [Checklists (Accessing Onsite Webpages From Offsite|offsite access)
    • Checklist for FCAL was discussed and suggestions made.
    • Sweep procedure will include checking on the lock on the FCAL Dark Room. Elton and accelerator personnel followed the sweep route and discussed how to include this yesterday afternoon. The accelerator procedure is being readied now for final review by Henry Robertson on Monday morning, who is currently away from the lab.
    • Checklist for BCAL has been completed up to the point of checking interlocks and taking LED data. The interlocks can be checked this evening in advance of power outage tomorrow morning. Elton asked for help taking the LED data this afternoon.
  6. FCAL HDFCAL log book
    • Matt noted that about 10 channels stopped working after the FCAL move. Checks of a couple of pmts showed no problems, so the cause of the dead channels is not yet known.
    • Currently the pmts are set using an ASCII file containing information about each base/pmt. In particular it contains parameters that use data taken during testing to determine the optimal voltage.
    • The voltages are set using information from this file to determine the set voltage.
    • Elton: Most of the pmt-based systems using the EPICs interface have a button to save the set voltages to a file, which can be restored at a later date. Manuel intends to implement this feature for the FCAL also
    • Matt: Does the system record which file was used to set the voltages? We do not think so.
  7. BCAL logbook
    • Mark put a study on attenuation lengths in the log book. Results are not understood.
    • Plot ratio of upstream to downstream pulse heights vs position determined by timing in the Flash ADCs
    • Data away from the bulk of data at the location of the trigger counters determines the attenuation length giving values around 1500 cm. This is a factor of 4 larger than previous measurements.
    • Andrei: Threshold effects may be important. Could use only events with large pulse heights. Also tail at about 20 ns may indicate this effect
    • Elton: Could be due to trigger bias, since hits away from the main peak probably are accidentals.
  8. Reconstruction and Simulation
    • Will M. has been looking at the simulations using commissioning target
    • Can now plot simple quantities with confidence
    • Non-pi0 and eta backgrounds are still large and preclude getting good numbers for rates in these channels.
    • Will L.: Willing to help with any problems in simulation or reconstruction. This is the first look at calorimetry data with the commissioning configuration, including a non-stardard target locations, etc. Also, it appears that the simulation contains many more noise hits that are not seen in the data. This could be because the data has thresholds of 5-6 mV not in the simulation. Still the simulation does not seem to be reproducing the data.
    • Will M. will send him plots and information to look at.
  9. Calibration
  10. Any other business